OpenStudy (anonymous):

Solve by the elimination method 0.3x-0.2y=4 0.5x+0.5y=-15/19

OpenStudy (saifoo.khan):

x = 7.368 y = -8.947

OpenStudy (anonymous):

That one was wrong! :(

OpenStudy (radar):

do you want the elimination process demonstrated?

OpenStudy (saifoo.khan):

the ques? :O

OpenStudy (anonymous):

no the answer.

OpenStudy (saifoo.khan):

wait, let me chk again.

OpenStudy (saifoo.khan):

im getting the same ans! :S

OpenStudy (saifoo.khan):

is the ques correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It tells you if the answer is correct or not.

OpenStudy (saifoo.khan):

:S

OpenStudy (radar):

Do you want to see the elimination method step by step?

OpenStudy (radar):

Or are you only interested in the answer?

OpenStudy (saifoo.khan):

how can u get this ans radat?

OpenStudy (saifoo.khan):

radar*

OpenStudy (radar):

saifoo.kahn I would like to repeat the steps and you critique the work.

OpenStudy (saifoo.khan):

i usd calc to find the answrs! :O

OpenStudy (radar):

Or anyone else. Here is what I did: Eq 1: .3x-0.2y=4 Eq 2: .5x+.5y =15/19 Step 1. Multiply Eq1 by 5 becomes: 1.5x-y=20 multiply Eq2 by 3 getting: 1.5x+1.5y=45/19 subtract the latter from the other 1.5x -y = 20 - 1.5x +1.5y=45/19 ---------------------- -2.5y =335/19 solve for y -2.5y 335 ------ = ----- 1 19 -2.5y (19) =335 -47.5y=335 y=-7052631579 oops I guessed I messed up on the sign the first time. Was their anything wrong with this effort?

OpenStudy (radar):

* put a decimal point between the 7 and 0 lol

OpenStudy (anonymous):

so what is x

OpenStudy (radar):

substitute the value in y and solve for x .5x + .5(-7.052631579)=15/19 .5x -3.526315789=.789473684 .5x=4.315789473 x=8.631578946 This again is different from the first solution for x both in sign and value.

OpenStudy (radar):

rachelgreen3, did you check my work?

OpenStudy (anonymous):

the answer was 140/19 -170/19

OpenStudy (saifoo.khan):

correct rachiel

OpenStudy (saifoo.khan):

rachelgreen3*

OpenStudy (saifoo.khan):

i said the same thing.

OpenStudy (saifoo.khan):

yes, i was CORRECT! :D

OpenStudy (radar):

I see now where I erred the second time, i copied the second equation wrong putting a 15/19 when it should of been a -15/19. Makes a big difference.

OpenStudy (saifoo.khan):

np Radar... Mistake happen.. (: